Find an equation of the line that passes through ​(−3​,−7​) and that is parallel to 4x+7y+14=0.
Give the answer in​ slope-intercept form.

Answer by Alan3354(69443)

Find an equation of the line that passes through ​(−3​,−7​) and that is parallel to 4x+7y+14=0.
Give the answer in​ slope-intercept form.
===================
Step 1, find the slope of 4x+7y+14=0
Call it m
--------------
Parallel lines have the same slope.
------
Step 2, use y-y1 = m*(x-x1) where (x1,y1) is the given point.